A cube of wood supporting 200gm mass just floats in water. When the mass is removed, the cube ruses by 2cm. What is the size of the cube?

Grade:12th pass

1 Answers

Aditi Chauhan
9 years ago

Let the edge of cube be l. When mass is on the cube of wood

200g + l3dwoodg = l 3waterg

? l3dwood = l 3water – 200 …..(i)

When the mass is removed

l3dwood = (l-2) l2dwater …..(ii)

From (i) and (ii)

l3dwater – 200 = (l-2) l2dwater

But dwater = 1

? l3 – 200 = l2 (l-2)

? l = 10cm